OpenStudy (anonymous):

On discovering that her family had a 70% risk of heart attack, Erin took a treadmill test to check her own potential of having a heart attack. The doctors told her that the reliability of the stress test is 67%. What is the probability that Erin will not have a heart attack and the test predicts that she will? A. 0.099 B. 0.201 C. 0.231 D. 0.469

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

@phi

OpenStudy (phi):

first, what is What is the probability that Erin will not have a heart attack ? use her family had a 70% risk of heart attack,

OpenStudy (phi):

what is the prob the test is correct ? multiply the two probabilities for the answer

OpenStudy (anonymous):

well all u hv to do is multiply the probabilities of % so 70/100 *67/100 = 7*67/1000 = 469/1000 = 0.469

OpenStudy (anonymous):

oh he was telling .-.

OpenStudy (anonymous):

its okay @aajugdar you did great too. I wish i could give you both metals cause i understood it both ways

OpenStudy (anonymous):

medal doesnt matter :P

OpenStudy (anonymous):

okayyy lml

OpenStudy (anonymous):

I got confused there but just in case the answer is C 0.231

OpenStudy (anonymous):

no answer C was wrong. Try 0.469 like stated above

OpenStudy (anonymous):

Thats wrong

OpenStudy (anonymous):

A. 0.099